# Rethinking Metrics and Benchmarks of Video Anomaly Detection

### v1

@@ Appendix D Details of Baselines
This paper evaluates existing models on large-scale datasets using our proposed metrics and benchmarks. To facilitate comprehensive comparison, we implement 10 state-of-the-art VAD methods, which can be categorized into four groups based on their technical characteristics: (1) traditional methods employing the common weakly-supervised paradigm without external knowledge; (2) text-enhanced methods that integrate vision-language pre-trained models for performance enhancement; (3) zero-shot methods leveraging large models’ generalization capabilities to mitigate data bias and perform VAD task without tuning; and (4) methods utilizing additional supervision, which improve accuracy at the expense of requiring labor-intensive annotation. It is worth noting that UR-DMU [5], GlanceVAD [32], and HolmesVAU [33] use the same backbone, thus allowing us to compare the performance of different supervision paradigms more fairly. For the experimental setup, we retrain RTFM [3] and MGFN [4] on UCF-Crime, retrain UR-DMU [5], PEL [9], and VadCLIP [10] on MSAD, and directly employ official open-source weights for all other methods.
Moreover, we observe significant discrepancies in existing implementations of AP computation. Specifically, some approaches calculate the interpolated area under the precision-recall curve (e.g., using sklearn.metrics.precision_recall_curve combined with sklearn.metrics.auc), while others adopt non-interpolated computation (e.g., employing sklearn.metrics.average_precision_score). However, as demonstrated in [51, 52], interpolation may yield overly optimistic AP estimations. To ensure fair comparison, we consistently implement the non-interpolated AP metric throughout our experiments.
Appendix E Further Evaluations
Visualizations of predictions of models with different LaAP.
As illustrated in Figure 3(a), PEL [9] demonstrates a significant improvement in prediction scores within the anomaly interval when compared to UR-DMU [5] and LAVAD [31]. Although LAVAD exhibits higher scores during the initial phase of the anomaly interval, UR-DMU manifests a lower false alarm rate in the normal part. From Figure 3(b), we could find that methods with higher LaAP (UR-DMU, RTFM, PEL) give higher scores within the anomaly interval, and the others (VadCLIP, MGFN) either have higher scores on the normal part or lower scores on the abnormal part. These visualization examples further demonstrate the reliability of the proposed LaAP in evaluating detection latency.
(a) An example from UCF-Crime.
(b) An example from MSAD.
Figure C: Examples of predictions of models with different LaAP. The dataset level LaAP for each model is listed in the legend, and note that LaAP, like AUC/AP, cannot be analyzed at the instance level.
Comparisons between AP and LaAP.
In Figure D, we visualize curves for AP and LaAP on two datasets. We find a large difference in UCF-Crime and a smaller difference in MSAD, possibly caused by the shorter video length on MSAD. Additionally, the curves for both LaAP and AP exhibit similar properties. Specifically: (1) They invariably traverse the critical $(0,1)$ coordinate point; (2) They do not necessarily traverse the $(1,0)$ point; and (3) Both demonstrate non-monotonically decreasing characteristics.
(a) Precision-LaRecall curve on UCF-Crime.
(b) Precision-Recall curve on UCF-Crime.
(c) Precision-LaRecall curve on MSAD.
(d) Precision-Recall curve on MSAD.
Figure D: Precision-LaRecall curve and Precision-Recall curve of existing models on UCF-Crime and MSAD.
Visualizations of FAR under different thresholds.
Figure E visually presents the FAR curves of all methods discussed in the main text, with specific annotation of FAR scores at both a balanced threshold (0.5) and a conservative threshold (0.8). Comparative analysis between performance on the original test set and the hard normal test set reveals significant elevation among the curves. This empirical evidence demonstrates that models exhibit systematically higher FAR across varying thresholds in the hard normal test set, thereby confirming the efficacy of our proposed benchmarks in revealing model overfitting to backgrounds. In addition, we can find that some models have higher FAR than others under the balanced threshold, but lower under the conservative threshold. By analyzing the curve, we can choose a more appropriate model according to the actual needs. For example, the user may select the threshold according to the requirement of FAR (e.g., $\leq 0.1\%$), and calculate the LaRecall for model selection.
(a) FAR curves on UCF-Crime (left) and UCF-HN (right).
(b) FAR curves on MSAD (left) and MSAD-HN (right).
Figure E: FAR curves of different methods at different thresholds, on the left are the results on original datasets (UCF-Crime and MSAD), and on the right are the results on the proposed hard normal datasets (UCF-HN and MSRA-HN).
Appendix F More Details of the Hard Normal Benchmarks
F.1 Construction of UCF-HN and MSAD-HN.
The purpose of constructing UCF-HN and MSAD-HN is to evaluate the degree of model overfitting by generating normal videos with the same scene for abnormal videos. We construct UCF-HN and MSAD-HN in three steps as follows:
  • Select Keyframes. To obtain high-quality images, we manually extracted key frames from anomaly videos in the training sets of UCF-Crime and MSAD. These keyframes met the following criteria: 1) No inclusion of abnormal events; 2) Absence of motion blur; 3) Minimal visible watermarks; 4) One keyframe per video.
  • Generate Videos. For UCF-HN, we employed the image-to-video generation mode of Vidu 2.0 model from vidu platform333www.vidu.cn/create/img2video with 1088×800 resolution at 32 fps. For MSAD-HN, we utilized the wan2.1_i2v_480p_14B model through ComfyUI444github.com/comfyanonymous/ComfyUI at 832×480 resolution with 16 fps. Each video was created by first generating a clip from the selected key frame, then producing an extended clip using the last frame of the first clip, followed by temporal concatenation. UCF-HN videos contain 15-second sequences while MSAD-HN videos comprise 10-second sequences. Manual prompt engineering ensured diverse event representations in generated videos. All key frames generated single videos except four key frames in UCF-HN, each producing five videos of different behaviors for granular analysis.
  • Recheck. During this phase, we re-generate videos exhibiting suboptimal quality while discarding frames with persistent synthesis artifacts. All retained videos were subjected to manual quality inspection to verify visual fidelity. This process culminated in the creation of two benchmark datasets: UCF-HN containing 100 videos, and MSAD-HN comprising 67 videos.
F.2 Comparisons with other VAD methods using synthetic data.
Figure F presents comparisons with existing synthetic data approaches in VAD. Notably, all compared methods differ in motivation from our work, with the comparison focused solely on generation quality. GTACrash [53] acquires crash data from gaming environments, exhibiting limited physical realism. UBNormal [27] employs 3D modeling software to produce animation, demonstrating suboptimal human motion naturalness and environmental integration. OVVAD [44] employs a text-to-image diffusion model coupled with an image animation model for frame generation, resulting in conspicuous visual artifacts and temporal inconsistencies. Astrid, M. et al. [54] generates anomalies via cut-paste operations, resulting in significant background inconsistency. Narwade, P. et al. [55] synthesizes alternative background states via image-to-image diffusion models while reinserting foreground objects through cut-paste operations, achieving natural motion patterns yet compromised foreground-background integration.
Our work pioneers the application of video diffusion models for benchmark generation, producing data with minimal artifacts, seamless human-environment integration, and unedited temporal continuity. The synthesized videos demonstrate superior realism compared to existing methods.
@@ Appendix F More Details of the Hard Normal Benchmarks
Figure F: Comparisons with other VAD methods using synthetic data [53, 27, 44, 54, 55]. Reference-based methods indicate methods that synthesize videos based on existing datasets, while reference-free methods do not rely on existing datasets.
F.3 More Visualizations of UCF-HN and MSAD-HN.
We provide more visualizations of our proposed synthetic benchmarks in Figures H,I,J.
F.4 Discussions of Video Surveillance Data Source.
Figure G: Analysis of existing VAD data sources in customizability, realism, scalability, and privacy/commercial concern.
Current VAD methods primarily focus on surveillance videos. We conduct a systematic discussion on data sources for surveillance video data as follows. The persistent challenge of data scarcity in VAD research is analyzed through four dimensions as in Figure G: Customizability, Realism, Scalability, and Privacy/Commercial Concern. Customizability refers to the variety of data you can get from this source; Realism refers to the similarity to real-world scenarios; Scalability refers to the difficulty of obtaining data at scale; Privacy/Commercial Concern addresses legal and commercial restrictions. These four dimensions are the key criteria to measure the quality of a data source.
Figure H: Examples of UCF-HN. (a) The original video contains an anomaly of a vehicle hitting an animal. The generated video shows vehicles and bicycles passing normally. (b) The original video contains an anomaly of an explosion. The generated video shows a cleaner walking around.
Real source denotes authentic surveillance footage containing actual anomaly events, which is typically inaccessible due to privacy constraints. Web source consists of publicly shared surveillance clips that enable large-scale data collection through crawling. While being the primary data source for many existing datasets, these videos often suffer from editing and watermarks. Researchers also find it hard to acquire high-quality normal videos, since uploaders generally do not share uninteresting normal surveillance footage. Movie source comprises films, TV series, etc., offering visually prominent anomalies (e.g., car crashes, explosions, gunfire) but also containing artistic embellishments that deviate from real-world scenarios. Recorded source involves staged anomaly simulations, providing controllable environments while inherently the presence of severe real-world anomalies. Game source leverages interactive virtual spaces and offers flexibility for data acquisition, but its physical simulations and graphical effects may demonstrate discrepancies from reality. 3D Animation source employs mature manual modeling techniques to create synthetic data, achieving higher visual fidelity than game engines by sacrificing real-time rendering capabilities. However, this method demands expertise and labor-intensive production processes. The emerging AIGC source utilizes generative models (e.g., diffusion models) with multimodal control mechanisms (e.g., text prompts, mask images, poses, depth maps). This method demonstrates unique advantages in generating various videos, especially sensitive content (e.g., NSFW, violent), while minimizing human intervention through an automated workflow. Recent advancements in this field promise continuous improvements in realism, customizability, and generation efficiency. Moreover, such synthetic data sources circumvent traditional privacy and commercial limitations.
In conclusion, AIGC represents a promising frontier for VAD research. This paper presents pioneering experiments utilizing AIGC, successfully addressing critical challenges unresolved by conventional data sources.
Appendix G Compute Resources
All experiments were performed on a single NVIDIA RTX 4090 GPU, including reproducing baselines, generating videos, and calculating metrics. We did not find an efficient calculation algorithm like AUC/AP as described in [56], however, the computation time on the existing data scale is controlled in the tens of seconds with a consumer-grade CPU.
Appendix H Dataset Statement
Re-annotations.
Our re-annotations of existing datasets are released exclusively for non-commercial academic purposes, containing no original video content and thereby eliminating potential privacy risks.
UCF-HN and MSAD-HN.
Our two synthetic benchmarks are based on existing UCF-Crime and MSAD datasets. The generated videos will be released exclusively for academic research or educational purposes. We did not associate the generated content with any real person.
Appendix I Potential Social Impact
This study advances the video anomaly detection field by providing a framework that facilitates the choice of more reliable and timely models, thereby enhancing their implementation in diverse security surveillance architectures. The proposed benchmarks are designed to evaluate if the model overfits on specific confounding elements (e.g., appearance, ethnic characteristics, or scenes), offering a privacy-agnostic evaluation protocol that enhances fairness and safety. This technology could potentially be misused for mass surveillance in combination with other identification technologies.
Figure I: Examples of MSAD-HN. (a) The original video contains a fire anomaly. The generated video shows a corgi coming into the room. (b) The original video contains an anomaly involving a glass door being broken. The generated video shows several teenagers talking to each other.
Figure J: Examples of UCF-HN where multiple videos are generated using the same keyframe. The original video contains a traffic accident anomaly. The generated videos comprise various normal events.

### v2

@@ V-A Baselines
This paper evaluates existing models on large-scale datasets using our proposed metrics and benchmarks. To facilitate comprehensive comparison, we implement 10 SOTA VAD methods, which can be categorized into four methodological groups: 1) traditional methods (RTFM [3], MGFN [4], UR-DMU [5]) employing the common weakly-supervised paradigm without external knowledge; 2) text-enhanced methods (PEL [9], VadCLIP [10], TEVAD [7], CLIP-TSA [8]) that integrate vision-language pre-trained models for performance enhancement; 3) zero-shot methods (LAVAD [34]) leveraging large models’ generalization capabilities to mitigate data bias and perform VAD task without tuning; and 4) methods utilizing additional supervision (GlanceVAD [35], HolmesVAU [36]), which improve accuracy at the expense of requiring labor-intensive annotation. It is worth noting that UR-DMU, GlanceVAD, and HolmesVAU use the same backbone, thus allowing us to compare the performance of different supervision paradigms more fairly. For the experimental setup, we retrain RTFM and MGFN on UCF-Crime, retrain UR-DMU, PEL, and VadCLIP on MSAD, and directly employ official open-source weights for all other methods.
V-B Benchmarking Existing Methods
We report the performance of SOTA VAD methods with our proposed probabilistic AUC/AP, LaAP and hard normal benchmarks in Tables IV,V. From Table IV, we observe that using probabilistic instead of original AUC/AP values significantly impacts method rankings, demonstrating that metric reliability is affected by annotator disagreements. This effect is most pronounced on UCF-Crime, while being relatively minor on XD-Violence and MSAD, which aligns with our observations of annotation discrepancy discussed in Section III-A. Furthermore, comparative analysis of LaAP metrics reveals that models with comparable AUC/AP may exhibit substantial differences in LaAP. For instance, on UCF-Crime, MGFN underperforms other methods in both AUC and AP metrics but achieves the highest LaAP among weakly-supervised approaches. Similarly on MSAD, UR-DMU and VadCLIP demonstrate identical AUC performance yet exhibit significant LaAP discrepancy. Experiments also confirm that methods incorporating additional supervisory signals achieve consistent improvements across all metrics.
From Table V, while most methods maintain low FAR ($\leq 10\%$) on the original test set, they suffer dramatic performance degradation on our proposed hard normal benchmarks. We observe a 42% average increase at 0.5 threshold, and 36% at 0.8 threshold, with some methods exceeding 70% FAR. This reveals severe scene overfitting in existing approaches. Comparative studies of supervision strategies using the same UR-DMU backbone reveal that clearer supervisory signals can mitigate overfitting risks. Although current anomaly videos in the training sets contain both normal and abnormal intervals, video-level weak supervision struggles to guide models in learning true anomaly patterns.
Figure 7: Consistency analysis between LaAP and AP, where each data point represents the metric values for a model at a specific training step.
Figure 8: Visualization of different score modification methods: Desc. re-arranges predictions within the abnormal area in descending order; Asc. does so in ascending order.
Figure 9: T-SNE visualization illustrating the distribution of reference abnormal videos, generated normal videos, and original normal videos.
V-C Reliability of LaAP
In Figure 9, we plot the AP and LaAP values and calculate the Spearman’s Rank-Order Correlation Coefficient (SROCC) during the training of VadCLIP, which shows strong correlation between LaAP and AP. To investigate the advantages of LaAP, we conduct a comparative analysis by modifying predicted scores from VadCLIP [10] and PEL [9] on UCF-Crime [16]. As illustrated in Figure 9, we simulate early/late detection through temporal reordering of anomaly scores within abnormal segments while preserving original score magnitudes. The Desc. makes the detection earlier by sorting scores in descending order, while Asc. achieves the opposite effect through ascending order. Table VI demonstrates LaAP’s latency awareness: While AUC/AP remain invariant, our LaAP effectively rewards earlier detections through its temporal weighting mechanism.
TABLE VI: Comparisons between AUC, AP, and LaAP under different score modification methods (visualized in Fig. 9). LaAP is able to effectively reward earlier detections.
Model Modification AUC AP LaAP
VadCLIP Desc. 0.900 0.486 0.783
Ori. 0.900 0.486 0.740
Asc. 0.900 0.486 0.675
PEL Desc. 0.886 0.439 0.744
Ori. 0.886 0.439 0.727
Asc. 0.886 0.439 0.645
V-D Domain Gap of UCF-HN and MSAD-HN
To validate the realism of our synthetic data, we compute Fréchet Inception Distance (FID) [48] and Fréchet Video Distance (FVD) [49] between UBNormal [30] and our method on real-world surveillance videos, which comprise videos of original test sets in UCF-Crime and MSAD. As in Table VII, significantly lower values demonstrate the superior realism of our benchmarks.
To further rule out the influence of the domain gap, we calculate the FAR on MSAD-HN using models trained on UCF-Crime, and compare it with results on the original UCF-Crime test set and UCF-HN. The videos in MSAD-HN are generated and contain scenes not present in the training set (i.e., UCF-Crime). If domain gap were the main cause of the observed false alarms, a high FAR on MSAD-HN would also be expected. However, experimental results in Tab. VIII demonstrate a relatively low false alarm rate on MSAD-HN, suggesting that the high false alarms are not attributable to the real-synthetic domain gap. This confirms that our benchmarks can reliably detect scene-specific overfitting.
(a) Comparison between a reference video and the corresponding generated video in UCF-HN.
@@ V-D Domain Gap of UCF-HN and MSAD-HN
(b) Comparisons with other VAD methods using synthetic data. Reference based methods indicate methods that synthesize videos based on existing datasets, while reference-free methods do not rely on existing datasets.
Figure 10: Comparison of synthetic video data generated by existing methods, including GTACrash [41], UBNormal [30], OVVAD [42], Astrid, M. et al. [43], and Narwade, P. et al. [44].
TABLE VII: Comparisons between UBNormal, UCF-HN and MSAD-HN with FID [48] and FVD [49] metrics.
Dataset FID $\downarrow$ FVD $\downarrow$
UBNormal [30] 202.6 698.0
UCF-HN 149.5 530.2
MSAD-HN 140.4 461.0
TABLE VIII: Analysis of domain gap with FAR0.5.
Test-set Model (trained on UCF-Crime)
RTFM [3] PEL [9] VadCLIP [10] MGFN [4]
UCF 0.129 0.005 0.024 0.033
UCF-HN 0.823 0.562 0.422 0.300
MSAD-HN 0.282 0.124 0.201 0.116
(a) Different annotation granularity.
(b) The start time is more consistent than the end time.
Figure 11: Examples of discrepancies across four rounds of annotations. Red regions indicate the annotated anomalies.
Figure 12: Example of predictions of models with different LaAP. The dataset-level LaAP for each model is listed in the legend.
We also provide visualized comparisons. Fig. 9 compares the similarity of the generated normal videos and the original normal videos with the abnormal videos in the feature space, where we use CLIP-ViT-B/16 [50] as the feature extractor and T-SNE to visualize sampled frames of each video. Our generated normal videos (hard negative samples) have closer proximity to abnormal videos than the original normal videos, which can better evaluate scene overfitting. In Fig. 10(a), we compares an original anomaly video (showing “Abuse”) with its generated normal counterpart, where violent actions are replaced by natural walking while maintaining perfect background coherence. Fig. 10(b) further compares with existing synthetic data approaches in VAD. Notably, all compared methods differ in motivation from our work, with the comparison focused solely on generation quality. Our method features more natural dynamics and a higher degree of foreground-background integration.
V-E Visualization
Fig. 11 presents two typical examples of annotation discrepancies, where they diverge in granularity and determination of ending (Finding 1). Fig. 12 visualizes different models’ predictions of a video, where PEL demonstrates a significant improvement in prediction scores within the anomaly interval when compared to UR-DMU and LAVAD. Although LAVAD exhibits higher scores during the initial phase of the anomaly interval, UR-DMU manifests a lower false alarm rate in the normal part. Note that LaAP, AUC, and AP do not have instance-level values due to their computational nature. Fig. 13 illustrates the prediction results for the original abnormal video and its generated normal counterpart. All three methods correctly raise high anomaly scores during for the crash video. However, they also raise high false alarms for the normal video that shares the same scene context. This phenomenon indicates that the models overfit to scene-specific cues rather than capturing the true anomaly semantics.
Figure 13: Examples of prediction results for the original abnormal video and its generated normal counterpart. All methods yield false alarms.
VI Conclusion
In this paper, we reveal critical limitations in current metrics and benchmarks of VAD through systematic analysis. To address these limitations, we propose new metrics for reliable and latency-sensitive evaluation (i.e., ProbAUC, ProbAP and LaAP). We also construct new benchmarks to evaluate the model generalizability (i.e., UCF-HN and MSAD-HN). We benchmark various SOTA VAD methods according to our new evaluation methods, providing novel perspectives on evaluating VAD algorithms.
Our findings suggest that generative data provides a promising solution to the scarcity of real-world data. Nevertheless, current methods remain inadequate in capturing complex motion patterns, preserving fine details, and ensuring temporal consistency. In future work, we aim to develop a scalable pipeline for generating high-quality video anomaly datasets.